# Fine-Grained Assertion-Based Test Selection

### v1

@@ Abstract
For large software applications, running the whole test suite after each code change is time- and resource-intensive. Regression test selection techniques aim at reducing test execution time by selecting only the tests that are affected by code changes. However, existing techniques select test entities at coarse granularity levels such as test class, which causes imprecise test selection and executing unaffected tests. We propose a novel approach that increases the selection precision by analyzing test code at statement level and treating test assertions as the unit for selection. We implement our fine-grained test selection approach in a tool called Selertion and evaluate it by comparing against two state-of-the-art test selection techniques using 11 open-source subjects. Our results show that Selertion increases selection precision for all the subjects. Our test selection reduces, on average, 63% of the overall test time, making regression testing up to 23% faster than the other techniques. Our results also indicate that subjects with longer test execution time benefit more by our fine-grained selection technique.
@@ I Introduction
Regression testing has been a vital process to ensure no existing functionality is broken when software evolves. However, even for large companies with allocated enormous resources to testing [1, 2, 3], it can be costly to rerun the whole test suite every time for the code changes. For example, Elbaum et al. [4] report that it costs seven weeks to execute the entire test suite for one of their industrial partner’s products. In addition to companies that conduct tests in continuous integration environments, individual developers often find themselves in a situation where they need to run regression tests multiple times a day before committing their changes to the codebase [5]. As a result, test suite optimization techniques that reduce test runtime such as regression test selection, minimization, and prioritization have become an inalienable part of software research and development [1, 6, 7]. Regression test selection (RTS) [8, 9] in particular aims to run only the tests that are affected by code changes. RTS techniques aim to be (1) safe, by selecting all tests affected by code changes, and (2) precise, by selecting fewer tests. Current techniques [10, 11, 9, 12] analyze code at different levels of granularity. Coarse-grained RTS techniques that analyze code at file-level [9] incur less analysis overhead but select more tests, i.e. they are less precise than finer-grained RTS techniques that analyze code at basic block or method level [10, 11]. However, existing RTS techniques typically focus on the granularity of production code (code under test), while select test entities at a relatively coarse granularity such as the test-class level. This can result in safe but imprecise test selection and, consequently, longer test execution times. The root cause of this issue lies in the fact that, in practice, only a portion of the test code within the selected test classes is impacted by code changes. For instance, when an atomic production method negate is changed in the Complex class from Apache Commons Math [13], the coarse-grained RTS techniques [12, 9] select the whole test class ComplexTest with 138 test cases. However, only seven out of the 138 test cases are truly affected by the changed method. Furthermore, empirical studies [14] have shown that it is a common practice for a single test case to contain multiple assertions. Therefore, only specific assertions within a test case might be influenced by code changes, leaving others unaffected. Our insight is that a finer-grained test code analysis can help to increase selection precision and accelerate test execution. As such, we focus on test methods and test assertions inside test methods, which are known to be correlated with a test suite’s fault detection ability [14]. We propose a novel test selection technique, called Selertion, which analyzes test entities at the fine-grained test assertion level. We treat each assertion inside a test method as the smallest unit of interest and slice the test method into executable fragments containing the targeted assertion and statements associated with it for selection. In comparison to coarse-grained RTS techniques, Selertion selects fewer tests more precisely, ultimately leading to a more efficient test execution process. Our work offers the following significant contributions:
@@
  • A method for statically slicing test methods into executable fragments for test each assertion.
@@
Our evaluation results reveal that Selertion selects 15.8% of tests, representing a precision improvement of 3–18% compared to existing state-of-the-art test selection techniques. Furthermore, Selertion significantly enhances efficiency by reducing, on average, 63% of the overall test execution time, resulting in a speed improvement of up to 23%.
@@ II Motivating Example
To illustrate our insight, we use real test cases from the ComplexTest class of Apache Commons Math [13] as a motivating example. Figure 1 shows two test methods testExp and testNegate that both have multiple assertions (i.e., Lines 5, 6, 8 in testExp and Lines 15, 16 in testNegate). When the production method Complex.negate is changed, all the 138 test cases in ComplexTest class will be selected if the selection is at test-class level. However, only seven (including the two shown in Figure 1 ) out of the 138 test cases are truly affected by the changed production method. Moreover, when focusing on the assertions in each test case, the first two assertions in testExp at Lines 5 and 6 neither check Complex.negate’s behavior nor have a dependency on it. Hence, it is not necessary to execute these two unaffected assertions. Meanwhile, if the two assertions in testNegate are assumed to fail because of the changes in Complex.negate, the execution will stop at the first failed assertion at Line 15, and the developer may overlook the second failed assertion at Line 16 when fixing the bug in Complex.negate. More time will be consumed by executing the test suite again to reveal the second failure at Line 16. The fine-grained test selection approach we propose is to divide a test method into smaller code fragments called assertion slices as the smallest units for selection. Each assertion slice includes the test assertion and previous statements associated with it. The assertion slices for a test method can be obtained by performing static intra-procedural backward slicing [15] on the variables of each assertion as the selected points of interest. Figure 2 presents the assertion slices for the two test methods testExp and testNegate in Figure 1 as an example. The grayed-out slices in Figure 2 are the ones that are selected when there are changes in the production method Complex.negate. Our fine-grained assertion-level selection technique selects only the assertions depending on the code changes and guarantees that all the selected assertions will be checked. Next we will present a more detailed description of our approach.
@@ III Approach
Figure 3 depicts an overview of our approach, which is composed of four main components. Change computation detects code changes at the class and method level in the production code. Assertion slicing computes test assertion slices for fine-grained selection by analyzing test code. Code instrumentation selectively instruments and executes both production and test code to collect dependencies. Finally, test selection selects test entities based on the information collected from the previous components. Using the terminology from related work [9, 12], the analysis (A) phase includes change computation, assertion slicing, and test selection; the collection (C) phase includes the code instrumentation and execution of instrumented subject; the execution (E) phase executes selected test entities.
@@ III-A Change Computation
Similar to state-of-the-art RTS tools, HyRTS [12] and Ekstazi [9], we detect code changes by comparing checksums. However, instead of tracing bytecode files, we compute checksums for source code files directly. We perform source code analysis because our fine-grained technique requires computation at statement level for test code to collect dependencies, and statement level analysis for bytecode will be indirect and may increase analysis complexity. The procedure to compute code changes is summarized in Algorithm 1. We trace code changes at multiple levels (i.e., method and class level). Given two versions of the code, we first compare the file-level checksums to detect the list of deleted, added, and changed files which are denoted by AF, DF, CF respectively at Line 1. Then for each file, we create class models to handle circumstances that one single file contains multiple classes (e.g., nested classes), although in most cases, there is one class per file (Line 2). We also handle enum types. Since enums can be treated as a special type of classes, we use class to represent both types in Algorithm 1. Intuitively, newly added classes (AC) and deleted classes (DC) are class-level changes, which are collected in the set $\delta^{c}$ at Line 3. Changed classes (CC) may include irrelevant changes that are not impacting the code, such as changes in comments, Javadoc, or indents. Such changes are excluded at the next step of the Algorithm (Line 4-12). For each changed class model, we separate the class body components into the following categories: class head (CH), methods (M) and other (OT). CH includes the class’s name, superclass or interfaces information. Methods (M) include all method declarations, including instance methods and constructors. Other (OT) includes other body declarations such as fields, except methods and classes. We store the smart checksums for each M, CH and OT separately. The reason to maintain a separate category OT is that, unlike bytecode computation in which field changes are reflected in method-level changes [12], we need to consider changed fields explicitly in source code analysis. To reduce the analysis overhead, we handle field changes all together and transfer them into a class-level change, because a field may be used by multiple methods. Therefore, for each changed class, if the OT or CH checksum varies, the corresponding class will be marked as changed (Line 5-6). If there are only method checksum changes without OT or CH change, the added (AM), deleted (DM) and changed methods (CM) will be stored as method-level changes in the set $\delta^{m}$ (Line 7-10). We also compute look up changes [10, 12] for added and deleted methods based on the inheritance information at Line 9 for the sake of safety in our analysis. In the initial run, because no previous checksums are available, every class is treated as a new class with class-level changes and the whole test suite is executed once.
@@ III-B Test Assertion Slicing
Let $S_{m}$ be the list of statements in the test method $m$. An assertion slice encompasses the assertion statement $asrt$, where $asrt\in S_{m}$ and its dependant statements $S_{m}[asrt]$, where $S_{m}[asrt]\subseteq S_{m}$. We use $\mathcal{A}$ to denote the total set of assertion slices for the test suite. $\mathcal{A}_{m}$ denotes a set of slices for the test method $m$. $\mathcal{A}_{m}[st]$ denotes a set of slices that contain the test statement $st$, where $st\in S_{m}$. An assertion $asrt$ is also a particular type of test statements.
@@
Using Figure 2 as an example, the test method ($m$) testExp’s assertion slices $\mathcal{A}_{m}$ are shown in the first three blocks. For the test statement $st$ at Line 3, $\mathcal{A}_{m}[st]$ contains one assertion slice shown in the first block. To obtain the assertion slices, we perform static intra-procedural backward slicing [15] for each test method based on the program dependence graph (PDG) using each assertion and its used variables as slicing criterion [16, 17, 18]. A PDG represents a procedure (i.e., the body of a test method) as a graph where nodes are statements or predicate expressions, and edges are control or data dependence. Generally speaking, test code has simpler structures than production code. Most test cases do not include conditionals such as if-else, for-loop, etc. To reduce the analysis overhead, we opted for only slicing test cases with explicit assertions and without conditional statements. Therefore, we apply data dependency analysis to create the PDG by capturing the variables used and defined for each test statement including the assertions. The global variables and methods annotated with @Before or @BeforeClass in test classes serve a purpose in points-to/alias analysis [17]; should always be included when executing the selected test entities, however, they do not need to be part of the data dependency analysis for individual test cases. Since we target an object-oriented language (Java), we pay more attention to object variable aliases in our analysis. For instance, when an object variable objA in the assertion statement is selected as the slicing criterion and its reference is passed to a method or constructor of another object variable objB, objA’s state may be changed by objB’s methods [17, 18]. If objA is passed to any field of objB, when objA is changed, objB’s state can also change. To make sure the slices do not miss any statements as data dependencies, we make a conservative assumption that method calls will change the states of the caller object and objects passed as arguments, which means a method invocation statement $st$ is associated with the targeted assertion $asrt$ if a variable in $asrt$ is used by $st$ as a caller object or passed to $st$ as an argument. Because assertions are a particular type of statement, if an assertion $asrt_{1}$ prior to the targeted assertion $asrt_{2}$ has a data dependency on $asrt_{2}$, $asrt_{1}$ will also be included in the slice for $asrt_{2}$.
@@ III-C Code Instrumentation
After obtaining the assertion slices, we collect test dependencies by capturing the method calls for each test statement. We instrument the production code to log the method signature for each production method invocation, and produce a separate set to keep track of method calls for each test statement inside test cases during test execution. For helper methods that are in test classes but are not test methods (i.e., without @Test annotation), we treat them the same way as production methods. Performing code instrumentation to collect dependencies at run-time for each new program revision will be inefficient. To reduce the overhead, we instrument each production and test class at initial run $V_{0}$, and store the instrumented subject as a separate copy (shown as the gray box in Figure 3). When a new revision $V_{i}$ is ready, we only need to re-instrument the changed files computed by Algorithm 1 and synchronize them to the instrumented copy of the subject, and execute the affected tests on the instrumented subject after test selection. Therefore, the dependency collection process is separately executed on the instrumented subject, which does not interfere with the execution of the original subject. We use $\mathcal{D}$ to denote all the test dependencies. For each test statement $st$, $\mathcal{D}(st)$ stands for a list of method invocations of $st$. Each method call is represented by its fully qualified method signature. Thus, it is easy to retrieve the class-level dependencies when needed. In practice, some test features might pose safety or efficiency challenges in tracing dependencies for each test statement. For parameterized test classes (annotated with @RunWith(Parameterized.class)), the method calls invoked by each test statement may vary for different inputs. It will increase analysis complexity to trace dependencies for each test statement with consideration of different inputs. We opted for collecting dependencies for each parameterized test class by logging all the method calls for different parameters together into one dependency file. When the dependency file contains any changed methods, all the test cases under different parameters in the parameterized test class are selected for execution. For the test classes that use inheritance, since each subclass may inherit test methods from the super class that are not explicitly available in the subclass, the approach to collect dependencies at test statement level is unsafe because the method-calls invoked by the inherited test methods will be missing for the subclass. Therefore, for the test classes that use inheritance (including both subclasses and superclasses), we trace dependencies at the test class level by logging method calls for the whole test class execution. Moreover, when a test class contains test methods calling other tests, it is hard to isolate each test method. For this case, we collect dependencies at the test class level as well. We identify those test classes that are parameterized, use inheritance or contain test methods calling other tests during static source code analysis for code instrumentation. Then, to trace dependencies at the test class level, instead of inserting code to log method calls for each test statement, we insert code only at the beginning and the end of each test class to keep track of the method calls. In addition to these specific test features, some test classes contain test methods that expect an exception to be thrown (e.g., annotated with @Test(expected=SomeException.class)). For those test methods, it is meaningless to trace method calls for each test statement, as the execution will be interrupted at a certain point and statements after that point will not be executed at all. Therefore, we collect dependencies for each test method instead. Moreover, in Section III-B, we choose to only slice the tests with explicit assertions and without conditional statements. For the test cases that contain conditional statements such as if-else or for-loop that are not sliced by the assertion slicing module, because they are not selected at fine-grained assertion level, it is sufficient to collect dependencies for each test method instead of each test statement to reduce overhead.

### v2

@@ Abstract
For large software applications, running the whole test suite after each code change is time- and resource-intensive. Regression test selection techniques aim at reducing test execution time by selecting only the tests that are affected by code changes. However, existing techniques select test entities at coarse granularity levels such as test class, which causes imprecise test selection and executing unaffected tests. We propose a novel approach that increases the selection precision by analyzing test code at statement level and treating test assertions as the unit for selection. We implement our fine-grained test selection approach in a tool called Selertion and evaluate it by comparing against two state-of-the-art test selection techniques using 11 open-source subjects. Our results show that Selertion increases selection precision for all the subjects. Our test selection reduces, on average, 63% of the overall test time, making regression testing 7–38% faster than the other techniques. Our results also indicate that subjects with longer test execution time benefit more by our fine-grained selection technique.
@@ I Introduction
Regression testing has been a vital process to ensure no existing functionality is broken when software evolves. However, even for large companies with allocated enormous resources to testing [1, 2, 3], it can be costly to rerun the whole test suite every time for the code changes. For example, Elbaum et al. [4] report that it costs seven weeks to execute the entire test suite for one of their industrial partner’s products. In addition to companies that conduct tests in continuous integration environments, individual developers often find themselves in a situation where they need to run regression tests multiple times a day before committing their changes to the codebase [5].
As a result, test suite optimization techniques that reduce test runtime such as regression test selection, minimization, and prioritization have become an inalienable part of software research and development [1, 6, 7]. Regression test selection (RTS) [8, 9] in particular aims to run only the tests that are affected by code changes. RTS techniques aim to be (1) safe, by selecting all tests affected by code changes, and (2) precise, by selecting fewer tests.
@@
Current techniques [10, 11, 9, 12] analyze code at different levels of granularity. Coarse-grained RTS techniques that analyze code at file-level [9] incur less analysis overhead but select more tests, i.e. they are less precise than finer-grained RTS techniques that analyze code at basic block or method level [10, 11]. However, existing RTS techniques typically focus on the granularity of production code (code under test), while select test entities at a relatively coarse granularity such as the test-class level. This can result in safe but imprecise test selection and, consequently, longer test execution times. The root cause of this issue lies in the fact that, in practice, only a portion of the test code within the selected test classes is impacted by code changes. For instance, when an atomic production method negate is changed in the Complex class from Apache Commons Math [13], the coarse-grained RTS techniques [12, 9] select the whole test class ComplexTest with 138 test cases. However, only seven out of the 138 test cases are truly affected by the changed method. Furthermore, empirical studies [14] have shown that it is a common practice for a single test case to contain multiple assertions. Therefore, only specific assertions within a test case might be influenced by code changes, leaving others unaffected.
Our insight is that a finer-grained test code analysis can help to increase selection precision and accelerate test execution. As such, we focus on test methods and test assertions inside test methods, which are known to be correlated with a test suite’s fault detection ability [14]. We propose a novel test selection technique, called Selertion, which analyzes test entities at the fine-grained test assertion level for objected-oriented Java programs. We treat each assertion inside a test method as the smallest unit of interest and slice the test method into executable fragments containing the targeted assertion and statements associated with it for selection. In comparison to coarse-grained RTS techniques, Selertion selects fewer tests more precisely, ultimately leading to a more efficient test execution process.
Our work offers the following significant contributions:
@@
  • A method for statically slicing test methods into executable fragments for testing each assertion.
@@
Our evaluation results reveal that Selertion selects 15.8% of tests, representing a precision improvement of 43%–53% compared to existing state-of-the-art test selection techniques. Furthermore, Selertion significantly enhances efficiency by reducing, on average, 63% of the overall test execution time, resulting in a speed improvement of 7–38%.
@@ II Motivating Example
To illustrate our insight, we use real test cases from the ComplexTest class of Apache Commons Math [13] as a motivating example. Figure 1 shows two test methods testExp and testNegate that both have multiple assertions (i.e., Lines 5, 6, 8 in testExp and Lines 15, 16 in testNegate). When the production method Complex.negate is changed, all the 138 test cases in ComplexTest class will be selected if the selection is at test-class level. However, only seven (including the two shown in Figure 1 ) out of the 138 test cases are truly affected by the changed production method. Moreover, when focusing on the assertions in each test case, the first two assertions in testExp at Lines 5 and 6 neither check Complex.negate’s behavior nor have a dependency on it. Hence, it is not necessary to execute these two unaffected assertions. Meanwhile, if the two assertions in testNegate are assumed to fail because of the changes in Complex.negate, the execution will stop at the first failed assertion at Line 15, and the developer may overlook the second failed assertion at Line 16 when fixing the bug in Complex.negate. More time will be consumed by executing the test suite again to reveal the second failure at Line 16.
The fine-grained test selection approach we propose is to divide a test method into smaller code fragments called assertion slices as the smallest units for selection. Each assertion slice includes the test assertion and previous statements associated with it. The assertion slices for a test method can be obtained by performing static intra-procedural backward slicing [15] on the variables of each assertion as the selected points of interest. Figure 2 presents the assertion slices for the two test methods testExp and testNegate in Figure 1 as an example. The grayed-out slices in Figure 2 are the ones that are selected when there are changes in the production method Complex.negate. Our fine-grained assertion-level selection technique selects only the assertions depending on the code changes and guarantees that all the selected assertions will be checked. Next we will present a more detailed description of our approach.
@@ III Approach
Figure 3 depicts an overview of our approach, which is composed of four main components. Change computation detects code changes at the class and method level in the production code. Assertion slicing computes test assertion slices for fine-grained selection by analyzing test code. Code instrumentation selectively instruments and executes both production and test code to collect dependencies. Finally, test selection selects test entities based on the information collected from the previous components.
Using the terminology from related work [9, 12], the analysis (A) phase includes change computation, assertion slicing, and test selection; the collection (C) phase includes the code instrumentation and execution of instrumented subject; the execution (E) phase executes selected test entities.
@@ III-A Change Computation
Similar to state-of-the-art RTS tools, HyRTS [12] and Ekstazi [9], we detect code changes by comparing checksums. However, instead of tracing bytecode files, we compute checksums for source code files directly. We perform source code analysis because our fine-grained technique requires computation at statement level for test code to collect dependencies, and statement level analysis for bytecode will be indirect and may increase analysis complexity.
The procedure to compute code changes is summarized in Algorithm 1. We trace code changes at multiple levels (i.e., method and class level). Given two versions of the code, we first compare the file-level checksums to detect the list of deleted, added, and changed files which are denoted by AF, DF, CF respectively at Line 1. Then for each file, we create class models to handle circumstances that one single file contains multiple classes (e.g., nested classes), although in most cases, there is one class per file (Line 2). We also handle enum types. Since enums can be treated as a special type of classes, we use class to represent both types in Algorithm 1. Intuitively, newly added classes (AC) and deleted classes (DC) are class-level changes, which are collected in the set $\delta^{c}$ at Line 3. Changed classes (CC) may include irrelevant changes that are not impacting the code, such as changes in comments, Javadoc, or indents. Such changes are excluded at the next step of the Algorithm (Line 4-12). For each changed class model, we separate the class body components into the following categories: class head (CH), methods (M) and other (OT). CH includes the class’s name, superclass or interfaces information. Methods (M) include all method declarations, including instance methods and constructors. Other (OT) includes other body declarations such as fields, except methods and classes. We store the smart checksums for each M, CH and OT separately.
The reason to maintain a separate category OT is that, unlike bytecode computation in which field changes are reflected in method-level changes [12], we need to consider changed fields explicitly in source code analysis. To reduce the analysis overhead, we handle field changes all together and transfer them into a class-level change, because a field may be used by multiple methods. Therefore, for each changed class, if the OT or CH checksum varies, the corresponding class will be marked as changed (Line 5-6). If there are only method checksum changes without OT or CH change, the added (AM), deleted (DM) and changed methods (CM) will be stored as method-level changes in the set $\delta^{m}$ (Line 7-10). We also compute look up changes [10, 12] for added and deleted methods based on the inheritance information at Line 9 for the sake of safety in our analysis. In the initial run, because no previous checksums are available, every class is treated as a new class with class-level changes and the whole test suite is executed once.
@@ III-B Test Assertion Slicing
Let $S_{m}$ be the list of statements in the test method $m$. An assertion slice encompasses the assertion statement $asrt$, where $asrt\in S_{m}$ and its dependant statements $S_{m}[asrt]$, where $S_{m}[asrt]\subseteq S_{m}$.
We use $\mathcal{A}$ to denote the total set of assertion slices for the test suite. $\mathcal{A}_{m}$ denotes a set of slices for the test method $m$. $\mathcal{A}_{m}[st]$ denotes a set of slices that contain the test statement $st$, where $st\in S_{m}$. An assertion $asrt$ is also a particular type of test statements.
@@
Using Figure 2 as an example, the test method ($m$) testExp’s assertion slices $\mathcal{A}_{m}$ are shown in the first three blocks. For the test statement $st$ at Line 3, $\mathcal{A}_{m}[st]$ contains one assertion slice shown in the first block. To obtain the assertion slices, we perform static intra-procedural backward slicing [15] for each test method based on the program dependence graph (PDG) using each assertion and its used variables as slicing criterion [16, 17, 18]. A PDG represents a procedure (i.e., the body of a test method) as a graph where nodes are statements or predicate expressions, and edges are control or data dependence. Generally speaking, test code has simpler structures than production code. Most test cases do not include conditionals such as if-else, for-loop, etc. To reduce the analysis overhead, we opted for only slicing test cases with explicit assertions and without conditional statements. Therefore, we apply data dependency analysis to create the PDG by capturing the variables used and defined for each test statement including the assertions. The global variables and methods annotated with @Before or @BeforeClass in test classes serve a purpose in points-to/alias analysis [17]; should always be included when executing the selected test entities, however, they do not need to be part of the data dependency analysis for individual test cases.
Since we target an object-oriented language (Java), we pay more attention to object variable aliases in our analysis. For instance, when an object variable objA in the assertion statement is selected as the slicing criterion and its reference is passed to a method or constructor of another object variable objB, objA’s state may be changed by objB’s methods [17, 18]. If objA is passed to any field of objB, when objA is changed, objB’s state can also change. To make sure the slices do not miss any statements as data dependencies, we make a conservative assumption that method calls will change the states of the caller object and objects passed as arguments, which means a method invocation statement $st$ is associated with the targeted assertion $asrt$ if a variable in $asrt$ is used by $st$ as a caller object or passed to $st$ as an argument. Because assertions are a particular type of statement, if an assertion $asrt_{1}$ prior to the targeted assertion $asrt_{2}$ has a data dependency on $asrt_{2}$, $asrt_{1}$ will also be included in the slice for $asrt_{2}$.
@@ III-C Code Instrumentation
After obtaining the assertion slices, we collect test dependencies by capturing the method calls for each test statement. We instrument the production code to log the method signature for each production method invocation, and produce a separate set to keep track of method calls for each test statement inside test cases during test execution. For helper methods that are in test classes but are not test methods (i.e., without @Test annotation), we treat them the same way as production methods.
Performing code instrumentation to collect dependencies at run-time for each new program revision will be inefficient. To reduce the overhead, we instrument each production and test class at initial run $V_{0}$, and store the instrumented subject as a separate copy (shown as the gray box in Figure 3). When a new revision $V_{i}$ is ready, we only need to re-instrument the changed files computed by Algorithm 1 and synchronize them to the instrumented copy of the subject, and execute the affected tests on the instrumented subject after test selection. Therefore, the dependency collection process is separately executed on the instrumented subject, which does not interfere with the execution of the original subject.
We use $\mathcal{D}$ to denote all the test dependencies. For each test statement $st$, $\mathcal{D}(st)$ stands for a list of method invocations of $st$. Each method call is represented by its fully qualified method signature. Thus, it is easy to retrieve the class-level dependencies when needed.
@@
In practice, some test features might pose safety or efficiency challenges in tracing dependencies for each test statement. For parameterized test classes (annotated with @RunWith(Parameterized.class)), the method calls invoked by each test statement may vary for different inputs. It will increase analysis complexity to trace dependencies for each test statement with consideration of different inputs. We opted for collecting dependencies for each parameterized test class by logging all the method calls for different parameters together into one dependency file. When the dependency file contains any changed methods, all the test cases under different parameters in the parameterized test class are selected for execution. For the test classes that use inheritance, since each subclass may inherit test methods from the super class that are not explicitly available in the subclass, the approach to collect dependencies at test statement level is unsafe because the method-calls invoked by the inherited test methods will be missing for the subclass. Therefore, for the test classes that use inheritance (including both subclasses and superclasses), we trace dependencies at the test class level by logging method calls for the whole test class execution. Moreover, when a test class contains test methods calling other tests, it is hard to isolate each test method. For this case, we collect dependencies at the test class level as well. We identify those test classes that are parameterized, use inheritance or contain test methods calling other tests during static source code analysis for code instrumentation. Then, to trace dependencies at the test class level, instead of inserting code to log method calls for each test statement, we insert code only at the beginning and the end of each test class to keep track of the method calls.
In addition to these specific test features, some test classes contain test methods that expect an exception to be thrown (e.g., annotated with @Test(expected=SomeException.class)). For those test methods, it is meaningless to trace method calls for each test statement, as the execution will be interrupted at a certain point and statements after that point will not be executed at all. Therefore, we collect dependencies for each test method instead. Moreover, in Section III-B, we choose to only slice the tests with explicit assertions and without conditional statements. For the test cases that contain conditional statements such as if-else or for-loop that are not sliced by the assertion slicing module, because they are not selected at fine-grained assertion level, it is sufficient to collect dependencies for each test method instead of each test statement to reduce overhead.